OpenStudy (anonymous):

[9.02] Where does the graph of y = 3x2 + 14x − 5 cross the x-axis? Medals to the best answer!

OpenStudy (anonymous):

3x^2 Sorry, that was a mistake

OpenStudy (anonymous):

crosses the x-axis when y = 0 3x^2 + 14x - 5 = 0 you should use the quadratic formula on the polynomial: ax^2 + bx + c above, which states that: \[x_{1,2} = \frac{ -b \pm \sqrt{b^{2} - 4ac} }{ ? }\] x1,2 is for the 2 x values that you will get (one for plus, one for minus). they can be the same

OpenStudy (anonymous):

rewriting the quadratic formula: \[x_{1,2} = \frac{ -b \pm \sqrt{b^{2} - 4ac}} { 2a }\]

OpenStudy (anonymous):

our case: \[x1,2 = \frac{ -14 \pm \sqrt{14^{2} - (4\times3\times (-5))} }{ 2(3) }\]

OpenStudy (anonymous):

How do I finish? @Euler271

OpenStudy (anonymous):

you have 2 answer: \[x1 = \frac{ -14 - \sqrt{256} }{ 6 }\] \[x2 = \frac{ -14 + \sqrt{256} }{ 6 }\]

OpenStudy (anonymous):

once you simplify with calculator, you are done

OpenStudy (anonymous):

Thank you!:)